the rectangular table for table tennis is x feet long and x-4 feet wide. write algebraic expressions for the perimeter and the area of the table. find the actual perimeter and area using x=9
Answer by fractalier(6550)

Perimeter P = 2L + 2W = 2x + 2(x-4) = 2x + 2x - 8 = 4x - 8
Area = LW = x(x - 4) = x^2 - 4x
If x = 9,
then P = 4(9) - 8 = 28 feet
and A = 9^2 - 4(9) = 45 feet^2